OpenStudy (twistrose):

What is the value of the expression 16 + 4 • 5 – 6 ÷ 2?

alones (alones):

16+(4)(5)−62 =16+20−62 =36−62 =36−3 =33

OpenStudy (twistrose):

thanks

alones (alones):

YOUR more than welcomed :)

OpenStudy (anonymous):

why did you bump this to the top

OpenStudy (twistrose):

on accident